OpenStudy (anonymous):

Find all the solutions for cosx cos(3x) - sinx sin(3x) = 0 ..on the interval [0,2pi]

OpenStudy (anonymous):

i think pi/4, 3pi/4, 5pi/4 and 7pi/4 :)

ganeshie8 (ganeshie8):

use cos(A+B) formula

OpenStudy (anonymous):

Thank you :3

OpenStudy (anonymous):

welcome :)

ganeshie8 (ganeshie8):

those are not the correct answers

OpenStudy (anonymous):

Would they be pi/8, 3pi/8, 5pi/8, 7pi/8, 9pi/8, 11pi/8, 13pi/8, 15pi/8?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

cos(4x) = 0 4x = pi/2 + npi x = pi/8 + npi/4 = pi/8, 3pi/8.... so yes !! good work !!!

OpenStudy (anonymous):

Thank you ><" So much

ganeshie8 (ganeshie8):

np :)